OpenStudy (anonymous):

1/4)^(2x+1) = 16 ^ (x+4)

OpenStudy (anonymous):

x=-9/4

jhonyy9 (jhonyy9):

write 1/4 = 2 on exponent (-2) and 16 = 2 on exonent 4

jhonyy9 (jhonyy9):

so and hence will get -2(2x+1)=4(x+4) -4x-2=4x+16 -18=8x x= -18/8 x= -9/4